Using dimensional analysis determine the volume of water in liters and gallons of a swimming pool with the

following dimensions: 50m L x 20m W x 2.5m deep.
Answer #1

Answer given dimension

dimensions: 50m L x 20m W x 2.5m deep.

we know volume = length * width* depth

volume = 50*20*2.5 m^3 = 2500 m^3

we know 1 m^3 = 1000 L

so that 2500 m^3 = 2500000 L

volume of pool = 2500000 L

and 1L = 0.264172 gallon

so that 2500000L = 2500000* 0.264172 = 660430 gallon

VOLUME OF POOL = 660430 gallon